# Universality Diagram of Phase Transitions in Long-range Statistical Systems

### v1

@@ III Long-range Simple Random Walk
A $d$-dimensional long-range simple random walk (LR-SRW), often referred to as a Lévy flight, is characterized by a jump probability distribution that decays algebraically with distance as $p(r)\sim r^{-(d+\sigma)}$ with $\sigma>0$ [bergersen1991, janssenLevyflightSpreadingEpidemic1999]. The transport dynamics are strictly governed by the exponent $\sigma$. Based on the convergence of the step distribution’s moments, the behavior can be classified into a standard diffusive regime and a complex superdiffusive regime, which is further subdivided as follows:
@@
  • Superdiffusive Regime: $\sigma\leq 2$. When $\sigma\leq 2$, the second moment diverges, and the walk is governed by Lévy stable distributions. In this regime, the transport is characterized by the typical displacement, $R_{\text{typ}}(t)\sim t^{1/\sigma}$, where the scaling exponent $\sigma$ varies across four distinct sub-regimes:
    • Marginal Superdiffusive: $\sigma=2$. At this boundary, the breakdown of the central limit theorem manifests as logarithmic corrections to the standard diffusion, leading to $R_{\text{typ}}(t)\sim(t\ln t)^{1/2}$ [bouchaud1990anomalous].
@@
    • Hyperballistic Regime: $\sigma<1$. The walk is dominated by extremely long-range jumps that effectively short-circuit the local lattice metric. The typical displacement grows super-linearly, $R_{\text{typ}}(t)\sim t^{1/\sigma}$ with $1/\sigma>1$, rendering the local details of the underlying lattice negligible.
@@ IV Long-range percolation model
Figure 1: Universality of the long-range bond percolation in the $(d,\sigma)$ plane. Six regimes are identified: the non-percolating region (I), the short-range universality regimes (II and V), and the long-range regimes (III–VI) with distinct scaling behaviors. The boundary at $\sigma=2$ separates the SR and LR universality classes, while $\sigma=1$ marks two different sectors in the nonclassical LR regimes.
@@
with $K\in(0,1]$ controlling the bond density, analogous to the reduced coupling in O($n$) spin models [grassberger_sir_2013, grassberger2013, liu_two-dimensional_2025]. For a given pair $(d,\sigma)$, a percolation threshold $K_{c}(d,\sigma)$ separates a non-percolating phase from a phase in which a cluster spans the system and diverges in size in the thermodynamic limit.
@@
In the $\sigma\to\infty$ limit, the model reduces to SR percolation, whose critical scaling behavior has been well established and is summarized in the previous section. Field-theoretically, SR percolation is governed by a $\phi^{3}$ theory with upper critical dimension $d_{c}=6$ [stauffer_introduction_2018, janssenLevyflightSpreadingEpidemic1999]. Below $d_{c}$, nontrivial fixed points control criticality. For $d\geq 2$, the percolation transition is continuous, while for $d=1$ the transition occurs only at $K_{c}=1$, corresponding to the absence of a nontrivial finite-threshold transition. By analogy with spin systems, we sometimes write $T_{c}=0$ to refer to the absence of a percolation transition. Above $d_{c}$, the Gaussian fixed point governs the critical exponents with $y_{t}^{\rm GF}=2$ and $y_{h}^{\rm GF}=\frac{2+d}{2}$, while periodic boundaries introduce an additional complete-graph asymptotics with $y_{t}^{\rm CG}=d/3$ and $y_{h}^{\rm CG}=2d/3$ [huang2018]. At $d=d_{c}$, these sectors coincide, and logarithmic corrections emerge in the finite-size scaling of both thermodynamic and geometric observables.
@@
Introducing LR bonds can drastically change the nature of the transition. For $d=1$, a nontrivial transition appears once the interaction becomes sufficiently long-ranged, i.e., when $\sigma\leq 1$, and at $\sigma=1$ the transition is of Berezinskii–Kosterlitz–Thouless (BKT) type [schulman1983, grassberger_sir_2013]. For $\sigma\leq d/3$, fluctuations are strongly suppressed and the LR Gaussian fixed point (LR-GFP) takes over. Dimensional analysis then yields mean-field-like exponents with $\nu=1/\sigma$ and $\eta=2-\sigma$. For intermediate values of $\sigma$, the interplay between the $k^{2}$ and $k^{\sigma}$ terms in the propagator determines whether criticality is SR- or LR-controlled. A crossover boundary at $\sigma_{*}=2-\eta_{\rm SR}$ was predicted by field-theoretical arguments [janssenLevyflightSpreadingEpidemic1999], whereas extensive numerical evidence in two-dimensional percolation and related O($n$) models consistently supports a sharp boundary at $\sigma_{*}=2$ [xiao_two-dimensional_2024, liu_two-dimensional_2025]. Complementary progress on the mathematical side shows that whenever $d>3\sigma$ and $\sigma<1$, the critical two-point function satisfies $G(r)\asymp r^{-d+\sigma}$, implying that the relation $\eta=2-\sigma$ is exact in this regime. More generally, this scaling is expected to hold throughout the region where an effective long-range condition remains valid. However, its extent near the SR–LR crossover, in particular for $1<\sigma<2$ and $d<3\sigma$, is still undetermined [hutchcroft_dimension_2025].
@@
The boundary of the no-transition $(T_{c}=0)$ regime can be intuitively inferred from the transport properties of LR-SRW. As discussed in Sec. III, the number of steps $t$ required to explore a typical length scale $R_{\rm typ}$ scales as $t\sim R_{\rm typ}^{2}$ for $\sigma>2$ and $t\sim R_{\rm typ}^{\sigma}$ for $\sigma<2$ [bouchaud1990anomalous]. To properly address the extensivity of an $d$-dimensional finite system, we consider a configuration where the total length of the random walk scales with the system volume, i.e., $O(L^{d})$. Thus, the establishment of LRO requires that the Lévy flights can effectively span the linear system size $L$ within $O(L^{d})$ steps. For the short-range regime ($\sigma>2$), traversing a distance $L$ requires $\sim L^{2}$ steps. Consequently, when $d<2$, the total number of steps $\sim L^{d}$ is insufficient to support such spanning trajectories, since $L^{d}\ll L^{2}$ in the thermodynamic limit, preventing the emergence of LRO. Conversely, in the long-range regime ($\sigma<2$), the superdiffusive transport implies that reaching distance $L$ requires only $\sim L^{\sigma}$ steps. For $d<\sigma$, the system lacks sufficient walk steps to bridge the distance $L$ in the thermodynamic limit, as $L^{d}\ll L^{\sigma}$. In this case, the correlations remain unable to span the system, again resulting in a disordered phase. This heuristic argument effectively delineates the boundaries of the no-transition regime at $\sigma=d$.
@@
Collectively, these field-theoretical, numerical, and rigorous insights delineate the topology of the universality diagram shown in Fig. 1, which contains six regimes: $T_{c}=0$, SR-LD, LR-LD-A, LR-LD-B, SR-HD, and LR-HD. Here, LD and HD denote low-dimensional and high-dimensional regimes, respectively. These regimes are separated by several structurally important boundaries that can be traced to (i) the existence of a nontrivial transition ($\sigma=d$ and $d=2$), (ii) the SR–LR crossover ($\sigma=2$), (iii) the relevance of interaction terms under coarse-graining ($\sigma=d/3$ and $d=6$), and (iv) a conjectured boundary at $\sigma=1$ that separates two types of LR regimes within the nonclassical regime.
@@
  • Regime I ($T_{c}=0$): $d<2$ and $\sigma>d$. Based on the LR SRW argument above, we conjecture that in this regime, the correlation length remains finite and no nontrivial transition exists for finite temperature $T>0$. Namely, the criticality is anomalous and at $T_{c}=0$, where all the lattice sites belong to a single giant cluster. In other words, the phase transition is first-order-like, with the thermal and magnetic renormalization exponents $y_{t}=y_{h}=d$. In particular, for $d=1$, both numerical and theoretical studies have demonstrated the absence of a phase transition for $\sigma>1$ [grassberger_sir_2013], which is consistent with this picture.
@@
For $d/3<\sigma\leq 2$ and $d<6$, long-range interactions modify the critical scaling, resulting in exponents that vary continuously with $\sigma$, characteristic of a nonclassical LR universality sector [fisher1972]. Within this sector, numerical and rigorous insights suggest a further subdivision into two distinct regimes:
  • Regime III (LR-LD-A): $1<\sigma\leq 2$ and $d/3<\sigma<d<6$. Here, the critical exponents vary continuously with $\sigma$. Along $\sigma=2$ in this range, the critical behavior remains LR-dominated, producing discontinuities of exponents and universal ratios as functions of $\sigma$ [liu_two-dimensional_2025]. High-precision numerical results indicate that both $y_{t}$ and $y_{h}$ deviate from the mean-field values of the LR Gaussian fixed point.
  • Regime IV (LR-LD-B): $d/3<\sigma<\min(1,d)$ for $d\in(0,3)$. Mathematical results show that $G(r)\asymp r^{-d+\sigma}$, i.e., $\eta=2-\sigma$, in this regime [hutchcroft_dimension_2025], and high-precision numerical simulations confirm this prediction [LRPercolation2]. At the same time, $y_{t}$ remains non-mean-field-like, for which no rigorous prediction is currently available.
@@
    so that macroscopic observables are governed by the CG scaling, whereas distance-dependent quantities probe the LR-GFP scaling sector. This two-scale FSS behavior in Regime VI has been confirmed numerically [LRPercolation2]. Furthermore, rigorous results show that the fractal dimension of critical clusters in this regime is $d_{\mathrm{F}}=2\sigma$ [hutchcroft_dimension_2025]. At the same time, the Fisher exponent remains $\tau=5/2$ even in the complete-graph model, and we therefore expect that this value persists in the LR-HD regime. Consequently, the geometric properties are expected to obey
@@
Beyond the six primary regimes, we further draw attention to several structurally important boundaries and special points in the universality diagram, where marginality or coincidence of scaling sectors gives rise to distinctive critical corrections:
@@
  • $(d,\sigma)=(1,1)$: The interaction is marginal and the transition is of BKT like, where the order parameter exhibits a discontinuous jump at $T_{c}$ [grassberger_sir_2013].
@@
It is worth emphasizing that the boundary $\sigma=1$ is conjectured based on the mathematical and numerical results. Mathematically, for $d/3<\sigma\leq 1$, it has been proved that $y_{h}=\frac{\sigma+d}{2}$ and $\eta=2-\sigma$ strictly. However, whether this result holds in the $\sigma>1$ regime remains an open question. High-precision Monte Carlo results show that the magnetic RG exponent is numerically consistent with the prediction $y_{h}=\frac{\sigma+d}{2}$ within error up to $\sigma\approx 4/3$ for $d=2$ [LRPercolation2]. Taken together, these findings suggest that $\sigma=1$ is at least a lower bound for a crossover between two qualitatively distinct LR-LD regimes.
Intriguingly, this boundary mirrors the kinetic transition of the LR-SRW at $\sigma=1$. As detailed in Sec. III, the transport dynamics shift from superdiffusive ($\sigma>1$) to hyperballistic ($\sigma\leq 1$), characterized by the typical displacement scaling $R_{\text{typ}}(t)\propto t^{1/\sigma}$. In the hyperballistic regime, the displacement grows superlinearly with time, i.e., $R_{\text{typ}}(t)\gg t$. This implies that a walker can span a spatial distance $R$ in fewer than $R$ time steps, signaling a fundamental change in the effective lattice connectivity and the bare propagator of the LR-GFP.
@@
Regime $(y_{t},\,y_{h})$
       SR
($y_{t}^{\rm SR},\,y_{h}^{\rm SR}$)
SR-GFP
(2, $\frac{d+2}{2}$)
LR-GFP
($\sigma$, $\frac{d+\sigma}{2}$)
CG
($\frac{d}{3},\frac{2d}{3}$)
SR-LD ($\checkmark$$\checkmark$) ($\times$$\times$) ($\times$$\times$) ($\times$$\times$)
LR-LD-A ($\times$$\times$) ($\times$$\times$) ($\times$$\times$) ($\times$$\times$)
LR-LD-B ($\times$$\times$) ($\times$$\times$) ($\times$$\checkmark$) ($\times$$\times$)
SR-HD ($\times$$\times$) ($\checkmark$$\checkmark$) ($\times$$\times$) ($\checkmark$$\checkmark$)
LR-HD ($\times$$\times$) ($\times$$\times$) ($\checkmark$$\checkmark$) ($\checkmark$$\checkmark$)
$T_{c}=0$ $(y_{t},y_{h})=(d,d)$
@@
LR-LD-B $\frac{d+\sigma}{2}$ $\frac{d+\sigma}{2}$ $\frac{2d}{d+\sigma}$
SR-HD $\frac{2}{3}d$ 44 $\frac{3}{2}$
LR-HD $\frac{2}{3}d$ $2\sigma$ $\frac{3}{2}$

### v2

@@ III Long-range Simple Random Walk
A $d$-dimensional long-range simple random walk (LR-SRW), often referred to as a Lévy flight, is characterized by a jump probability distribution that decays algebraically with distance as $p(r)\sim r^{-(d+\sigma)}$ with $\sigma>0$ [5, 31]. The transport dynamics are strictly governed by the exponent $\sigma$. Based on the convergence of the step distribution’s moments, the behavior can be classified into a standard diffusive regime and a complex superdiffusive regime, which is further subdivided as follows:
@@
  • Superdiffusive Regime: $0<\sigma\leq 2$. When $\sigma\leq 2$, the second moment diverges, and the walk is governed by Lévy stable distributions. In this regime, the transport is characterized by the typical displacement, $R_{\text{typ}}(t)\sim t^{1/\sigma}$, where the scaling exponent $\sigma$ varies across four distinct sub-regimes:
    • Marginal Superdiffusive: $\sigma=2$. At this boundary, the breakdown of the central limit theorem manifests as logarithmic corrections to the standard diffusion, leading to $R_{\text{typ}}(t)\sim(t\ln t)^{1/2}$ [7].
@@
    • Hyperballistic Regime: $0<\sigma<1$. The walk is dominated by extremely long-range jumps that effectively short-circuit the local lattice metric. The typical displacement grows super-linearly, $R_{\text{typ}}(t)\sim t^{1/\sigma}$ with $1/\sigma>1$, rendering the local details of the underlying lattice negligible.
@@ IV Long-range percolation model
Figure 1: Universality of the long-range bond percolation in the $(d,\sigma)$ plane. Seven regimes are identified: the non-percolating region (I), the short-range universality regimes (II and V), the long-range regimes (III–VI), and the complete-graph-like region (CG, VII) with distinct scaling behaviors. The boundary at $\sigma=2$ separates the SR and LR universality classes, while $\sigma=1$ marks two different sectors in the nonclassical LR regimes.
@@
with $K\in(0,1]$ controlling the bond density, analogous to the reduced coupling in O($n$) spin models [21, 22, 38]. For a given pair $(d,\sigma)$, a percolation threshold $K_{c}(d,\sigma)$ separates a non-percolating phase from a phase in which a cluster spans the system and diverges in size in the thermodynamic limit.
@@
In the $\sigma\to\infty$ limit, the model reduces to SR percolation, whose critical scaling behavior has been well established and is summarized in the previous section. Field-theoretically, SR percolation is governed by a $\phi^{3}$ theory with upper critical dimension $d_{c}=6$ [48, 31]. Below $d_{c}$, nontrivial fixed points control criticality. For $d\geq 2$, the percolation transition is continuous, while for $d=1$ the transition occurs only at $K_{c}=1$, corresponding to the absence of a nontrivial finite-threshold transition. By analogy with spin systems, we sometimes write $T_{c}=0$ to refer to the absence of a percolation transition. Above $d_{c}$, the Gaussian fixed point governs the critical exponents with $y_{t}^{\rm GF}=2$ and $y_{h}^{\rm GF}=\frac{2+d}{2}$, while periodic boundaries introduce an additional complete-graph asymptotics with $y_{t}^{\rm CG}=d/3$ and $y_{h}^{\rm CG}=2d/3$ [26]. At $d=d_{c}$, these sectors coincide, and logarithmic corrections emerge in the finite-size scaling of both thermodynamic and geometric observables.
@@
Introducing LR bonds can drastically change the nature of the transition. For $0<\sigma\leq d/3$, fluctuations are strongly suppressed and the LR Gaussian fixed point (LR-GFP) takes over. Dimensional analysis then yields mean-field-like exponents with $\nu=1/\sigma$ and $\eta=2-\sigma$. For intermediate values of $\sigma$, the interplay between the $k^{2}$ and $k^{\sigma}$ terms in the propagator determines whether criticality is SR- or LR-controlled. A crossover boundary at $\sigma_{*}=2-\eta_{\rm SR}$ was predicted by field-theoretical arguments [31], whereas extensive numerical evidence in two-dimensional percolation and related O($n$) models consistently supports a sharp boundary at $\sigma_{*}=2$ [56, 38]. Complementary progress on the mathematical side shows that whenever $d>3\sigma$ and $\sigma<1$, the critical two-point function satisfies $G(r)\asymp r^{\sigma-d}$, implying that the relation $\eta=2-\sigma$ is exact in this regime. More generally, this scaling is expected to hold throughout the region where an effective long-range condition remains valid. However, its extent near the SR–LR crossover, in particular for $1<\sigma<2$ and $d<3\sigma$, is still undetermined [30].
@@
The boundary of the no-transition $(T_{c}=0)$ regime can be intuitively inferred from the transport properties of LR-SRW. As discussed in Sec. III, the number of steps $t$ required to explore a typical length scale $R_{\rm typ}$ scales as $t\sim R_{\rm typ}^{2}$ for $\sigma>2$ and $t\sim R_{\rm typ}^{\sigma}$ for $0<\sigma<2$ [7]. To properly address the extensivity of an $d$-dimensional finite system, we consider a configuration where the total length of the random walk scales with the system volume, i.e., $O(L^{d})$. Thus, the establishment of LRO requires that the Lévy flights can effectively span the linear system size $L$ within $O(L^{d})$ steps. For the short-range regime ($\sigma>2$), traversing a distance $L$ requires $\sim L^{2}$ steps. Consequently, when $d<2$, the total number of steps $\sim L^{d}$ is insufficient to support such spanning trajectories, since $L^{d}\ll L^{2}$ in the thermodynamic limit, preventing the emergence of LRO. Conversely, in the long-range regime ($0<\sigma<2$), the superdiffusive transport implies that reaching distance $L$ requires only $\sim L^{\sigma}$ steps. For $d<\sigma$, the system lacks sufficient walk steps to bridge the distance $L$ in the thermodynamic limit, as $L^{d}\ll L^{\sigma}$. In this case, the correlations remain unable to span the system, again resulting in a disordered phase. This heuristic argument effectively delineates the boundaries of the no-transition regime at $\sigma=d$.
@@
Collectively, these field-theoretical, numerical, and rigorous insights delineate the topology of the universality diagram shown in Fig. 1, which contains seven regimes: $T_{c}=0$, SR-LD, LR-LD-A, LR-LD-B, SR-HD, LR-HD, and CG-like. Here, LD and HD denote low-dimensional and high-dimensional regimes, respectively. These regimes are separated by several structurally important boundaries that can be traced to (i) the existence of a nontrivial transition ($\sigma=d$ and $d=2$), (ii) the SR–LR crossover ($\sigma=2$), (iii) the relevance of interaction terms under coarse-graining ($\sigma=d/3$ and $d=6$), and (iv) a conjectured boundary at $\sigma=1$ that separates two types of LR regimes within the nonclassical regime.
@@
  • Regime I ($T_{c}=0$): $d<2$ and $\sigma>d$. Based on the LR SRW argument above, we conjecture that in this regime, the correlation length remains finite and no nontrivial transition exists for finite temperature $T>0$. Namely, the criticality is anomalous and at $T_{c}=0$, where all the lattice sites belong to a single giant cluster. In other words, the phase transition is first-order-like, with the thermal and magnetic renormalization exponents $y_{t}=y_{h}=d$. In particular, for $d=1$, both numerical and theoretical studies have demonstrated the absence of a phase transition for $\sigma>1$ [21], which is consistent with this picture.
@@
For $d/3<\sigma\leq 2$ and $d<6$, long-range interactions modify the critical scaling, resulting in exponents that vary continuously with $\sigma$, characteristic of a nonclassical LR universality sector [19]. Within this sector, numerical and rigorous insights suggest a further subdivision into two distinct regimes:
  • Regime III (LR-LD-A): $1<\sigma\leq 2$ and $d/3<\sigma<d<6$. Here, the critical exponents vary continuously with $\sigma$. Along $\sigma=2$ in this range, the critical behavior remains LR-dominated, producing discontinuities of exponents and universal ratios as functions of $\sigma$ [38]. High-precision numerical results indicate that both $y_{t}$ and $y_{h}$ deviate from the mean-field values of the LR Gaussian fixed point.
  • Regime IV (LR-LD-B): $d/3<\sigma<\min(1,d)$ for $d\in(0,3)$. Mathematical results show that $G(r)\asymp r^{-d+\sigma}$, i.e., $\eta=2-\sigma$, in this regime [30], and high-precision numerical simulations confirm this prediction [37]. At the same time, $y_{t}$ remains non-mean-field-like, for which no rigorous prediction is currently available.
    For $d=1$ and $\sigma=1$ the transition is of Berezinskii–Kosterlitz–Thouless (BKT) type [47, 21].
@@
    so that macroscopic observables are governed by the CG scaling, whereas distance-dependent quantities probe the LR-GFP scaling sector. This two-scale FSS behavior in Regime VI has been confirmed numerically [37]. Furthermore, rigorous results show that the fractal dimension of critical clusters in this regime is $d_{\mathrm{F}}=2\sigma$ [30]. At the same time, the Fisher exponent remains $\tau=5/2$ even in the complete-graph model, and we therefore expect that this value persists in the LR-HD regime. Consequently, the geometric properties are expected to obey
@@
Additionally, Regime VII (CG): $d\leq\sigma\leq 0$. Due to the divergence of the integration $\int r^{-(d+\sigma)}r^{d-1}dr$, one has to first consider a finite system of side length $L$ and then introduce a rescaled factor $L^{-\sigma}$ such that the bond occupation probability becomes $p(r)=(1/L^{\sigma})K/r^{d+\sigma}$, satisfying the extensivity of the system. When $\sigma=-d$, the interaction becomes distance-independent, and the lattice reduces to a complete graph (CG). The spatial fluctuations are so suppressed that the thermal RG exponent associated with the LR Gaussian fixed point, $y_{t}^{\rm GF}=\sigma<0$, becomes irrelevant, and the critical behaviors are solely described by the CG-asymptotics, with critical exponents $(y_{t}^{\rm CG},y_{h}^{\rm CG})=(d/3,2d/3)$.
Beyond the seven primary regimes, we further draw attention to several structurally important boundaries and special points in the universality diagram, where marginality or coincidence of scaling sectors gives rise to distinctive critical corrections:
@@
  • $(d,\sigma)=(1,1)$: The interaction is marginal and the transition is of BKT like, where the order parameter exhibits a discontinuous jump at $T_{c}$ [21].
@@
It is worth emphasizing that the boundary $\sigma=1$ is conjectured based on the mathematical and numerical results. Mathematically, for $d/3<\sigma\leq 1$, it has been proved that $y_{h}=\frac{\sigma+d}{2}$ and $\eta=2-\sigma$ strictly. However, whether this result holds in the $\sigma>1$ regime remains an open question. High-precision Monte Carlo results show that the magnetic RG exponent is numerically consistent with the prediction $y_{h}=\frac{\sigma+d}{2}$ within error up to $\sigma\approx 4/3$ for $d=2$ [37]. Taken together, these findings suggest that $\sigma=1$ is at least a lower bound for a crossover between two qualitatively distinct LR-LD regimes.
Intriguingly, this boundary mirrors the kinetic transition of the LR-SRW at $\sigma=1$. As detailed in Sec. III, the transport dynamics shift from superdiffusive ($\sigma>1$) to hyperballistic ($0<\sigma\leq 1$), characterized by the typical displacement scaling $R_{\text{typ}}(t)\propto t^{1/\sigma}$. In the hyperballistic regime, the displacement grows superlinearly with time, i.e., $R_{\text{typ}}(t)\gg t$. This implies that a walker can span a spatial distance $R$ in fewer than $R$ time steps, signaling a fundamental change in the effective lattice connectivity and the bare propagator of the LR-GFP.
@@
Regime $(y_{t},\,y_{h})$        SR ($y_{t}^{\rm SR},\,y_{h}^{\rm SR}$) SR-GFP (2, $\frac{d+2}{2}$) LR-GFP ($\sigma$, $\frac{d+\sigma}{2}$) CG ($\frac{d}{3},\frac{2d}{3}$)
SR-LD     (II) ($\checkmark$$\checkmark$) ($\times$$\times$) ($\times$$\times$) ($\times$$\times$)
LR-LD-A (III) ($\times$$\times$) ($\times$$\times$) ($\times$$\times$) ($\times$$\times$)
LR-LD-B (IV) ($\times$$\times$) ($\times$$\times$) ($\times$$\checkmark$) ($\times$$\times$)
SR-HD     (V) ($\times$$\times$) ($\checkmark$$\checkmark$) ($\times$$\times$) ($\checkmark$$\checkmark$)
LR-HD     (VI) ($\times$$\times$) ($\times$$\times$) ($\checkmark$$\checkmark$) ($\checkmark$$\checkmark$)
CG           (VII) ($\times$$\times$) ($\times$$\times$) ($\times$$\times$) ($\checkmark$$\checkmark$)
$T_{c}=0$      (I) $(y_{t},y_{h})=(d,d)$
@@
LR-LD-B  (IV) $\frac{d+\sigma}{2}$ $\frac{d+\sigma}{2}$ $\frac{2d}{d+\sigma}$
SR-HD     (V) $\frac{2}{3}d$ 44 $\frac{3}{2}$
LR-HD     (VI) $\frac{2}{3}d$ $2\sigma$ $\frac{3}{2}$